Fermi National Laboratory


Accelerator Update for November 2, 2005 - November 4, 2005

Wednesday November 2
The day shift began with Operations monitoring store 4480 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with NuMI. and MiniBooNE taking beam.

The Meson MCenter experiment resumed taking beam at 8:15 AM.

At 8:37 AM, Operations shut off the Main Injector (MI) beam so that experts could conduct RF studies. MI resumed sending beam at 9:17 AM.

At 12:02 PM, Operations shut off the MI beam so that experts could switch their RF back to its original configuration. MI resumed normal operations at 12:35 PM.

At 2 PM, a Tevatron (TeV) RF station (TRF1) tripped off.

Operations began preparing the Recycler at 2 PM to stash antiprotons. The stash ended and Pbar resumed stacking at 3:02 PM.

MCenter resumed taking beam at 4:05 PM.

Operations began preparing the Recycler at 6:49 PM to stash antiprotons. The stash ended and Pbar resumed stacking at 7:30 PM.

MCenter resumed taking beam at 7:39 PM.

Store 4480 aborted at 8:20 PM when TRF6 tripped off. Operations contacted RF experts.

Operators began preparing CDF and D0 for access at 8:56 PM. TeV experts accessed TeV sector A4 to repair a spool piece.

Operations began preparing the Recycler at 11:14 PM to stash antiprotons.

CDF and D0 came out of access at 11:21 PM.

Thursday November 3
The midnight shift began with the Recycler stashing, with Pbar and MCenter off due to the stash, and with MiniBooNE and NuMI taking beam.

The stash ended and Pbar resumed stacking at 12:09 AM.

Operations put the TeV through a dry squeeze at 1:15 AM.

Operations put the TeV into shot setup at 1:39 AM.

At 2:44 AM, Operations reported a P1 transfer line vacuum burst. Operators had the vacuum recovered within five minutes.

At 4:08 AM, TRF1 tripped off.

Operations established store 4481 at 4:11 AM with an initial luminosity of 148.83E30.

Operations began preparing the Recycler at 4:32 AM to stash. The stash ended and Pbar resumed stacking at 5:23 AM. MCenter also resumed taking beam.

Operations shut off MiniBooNE at 8:02 AM for an access.

Operations began preparing the Recycler at 8:36 AM to stash antiprotons.

At 8:59 AM, the MI power supply beam permit dropped out. Operators had the beam line back up with in five minutes.

The stash ended and Pbar resumed stacking at 9:15 AM. MCenter also resumed taking beam.

Operations reestablished beam to MiniBooNE at 11:08 AM.

Operations began preparing the Recycler at 1:27 PM to stash. The stash ended and Pbar and MCenter resumed normal operations by 2:21 PM.

At 5:45 PM, Operations reported that they had lost the Switchyard (SY) beam permit due to a power supply (HP3DS) overheating. Operators had to wait until 6:36 PM before the supply cooled off enough to reset. MCenter resumed taking beam.

Operations began preparing the Recycler at 11:35 PM to stash.

Friday November 4
The midnight shift began with Operations monitoring store 4481 and the stash, with Pbar and MCenter off due to the Recycler stash, and with MiniBooNE and NuMI taking beam.

The stash ended and Pbar and MCenter resumed normal operations around 12:30 AM.

Operations began preparing the Recycler at 3:55 AM to stash. The stash ended and Pbar and MCenter resumed normal operations by 4:51 AM.

TeV experts began a helix end-of-store study around 8 AM.

The Plans for Friday
The plans for today are to stack, stash, and establish stores on a 24-hour cycle.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/4/2005   email Fermilab